# Applied-Field Magnetoplasmadynamic Thrusters for Deep Space Exploration

### v1

@@ 2.2.5 Onset Phenomena
According to this theory, the onset occurs when the current density at the anode reaches a level where there is not enough thermal energy to sustain the plasma discharge, which leads to the detachment of the plasma from the anode. The anode becomes “starved” of electrons, disrupting the thruster’s stable operation.
@@ 3.3 Thermal Control
Many companies have bespoke patented thermal management technologies with little literature available in the public domain, but much of these developments pertain to the material design of the thermal protection system and unique thermal properties of these materials being exploited in order to serve as effective insulation between the thruster and the HTS coil within the Applied-Field module. Thermal control systems in AF-MPDTs are very similar to thermal control systems used for atmospheric reentry on chemical rockets.
@@ 3.3.2 Active Methods
Film cooling involves the injection of a coolant fluid along the vehicle’s surface, creating a protective film that protects the structure from extreme heat. Film cooling is currently used in high-speed aircraft and rocket engines. Computer simulations and wind tunnel tests continue to improve the effectiveness of film cooling in various flight conditions [32]. Research into the use of microjet arrays and advanced materials could enhance film cooling’s effectiveness for future deep space missions [33].
@@ 3.4.3 Argon
Argon has been another common choice for AF-MPDT propellants due to its availability and simplicity of handling. However, argon typically results in lower performance metrics compared to lithium and hydrogen [38]. Efficiencies for argon propellants in the 10-100 kW range are generally in the 5-25% range, with specific impulses between 700-1800s [37]. Despite its lower performance, argon is still valuable for certain applications due to its operational stability and ease of use.
@@ 4 State of the Art: AF-MPDTs
Current AF-MPDT research and development is focused on the use of HTS coils to generate the magnetic field. Generally, previously built AF-MPDTs have had a magnetic field strength of up to 0.4 Teslas [39]. Creating a magnetic field with a magnetic field strength over 1 Tesla can theoretically maximize the performance of the thruster[19]. However, with conventional electromagnets, that strength of a magnetic field requires bulky designs with high power usage. The high current densities of superconductors enable the creation of lighter, more compact, electromagnets, allowing the thruster to avoid the heavy mass and power consumption tolls of the conventional electromagnet[40]. Multiple entities around the world are currently engaged with superconducting AF-MPDT research, including groups in Russia, Germany, China, and New Zealand[19].
@@ 4.1 New Zealand - Robinson Research Institute (RRI)
4.2 Germany - Neutron Star Systems (NSS)
Neutron Star Systems is a German startup that has built a 5 kW AF-MPDT and is currently developing a 100 kW AF-MPDT. The 5 kW AF-MPDT features a light and compact superconductor and cryocooler system. They are currently researching coil configurations, cryogenic technologies, and magnetic field shapes. Utilizing machine learning, NSS is using 3D electromagnet models to optimize the field strength for high thrust performance[19].
@@ 4.4 China - Chinese Academy of Sciences (CAS)
In 2022, the Chinese Academy of Sciences successfully designed and tested a 150 kW AF-MPDT with superconducting coils. They reported an efficiency of 76% and a specific impulse of over 5600 seconds. The thruster also demonstrated and confirmed the long-held theory that more powerful magnetic fields can reduce cathode erosion and offer a solution to AF-MPDT lifetime issues[39][19].
@@ Acknowledgements
The authors would like to thank CCIR for funding support and the opportunity to conduct research on this topic.

### v2

@@ 2.2.5 Onset Phenomena
According to this theory, the onset occurs when the current density at the anode reaches a level where there is not enough thermal energy to sustain the plasma discharge, which leads to the detachment of the plasma from the anode. The anode becomes ”starved” of electrons, disrupting the thruster’s stable operation.
@@
Recent findings have appeared to strongly support the explanation of onset phenomena through the Anode Starvation Theory. Notably, Professor Heiermann’s numerical simulations at the University of Stuttgart affirm this effect. It shows that as the density of ions decreases in front of the anode, the resulting ”starvation” leads to thruster instabilities under high electric currents. As a result, energy transfer becomes inefficient, and increased wear on the anode is observed, further disrupting stable operation[16].
@@ 3.3 Thermal Control
Many companies have bespoke patented thermal management technologies with little literature available in the public domain, but much of these developments pertain to the material design of the thermal protection system and unique thermal properties of these materials being exploited in order to serve as effective insulation between the thruster and the HTS coil within the Applied-Field module. Thermal control systems in AF-MPDTs bear similarity to heat shield technologies used in atmospheric entry as well as to the thermal shielding applied in chemical rockets.
@@ 3.3.2 Active Methods
Film cooling involves the injection of a coolant fluid along the vehicle’s surface, creating a protective film that protects the structure from extreme heat. Film cooling is currently used in high-speed aircraft and rocket engines. Computer simulations and plasma wind tunnel tests can be used for research purposes to continue to improve the effectiveness of film cooling [33]. Research into the use of microjet arrays and advanced materials could enhance film cooling’s effectiveness for future deep space missions [34].
@@ 3.4.3 Argon
Argon has been another common choice for AF-MPDT propellants due to its availability and simplicity of handling. Argon typically results in lower performance metrics compared to lithium and hydrogen [39]. Efficiencies for argon propellants in the 10-100 kW range are generally in the 5-25% range, with specific impulses between 700-1800s [38].
However, radial segmentation of the mass flow rate between the cathode and anode has been found to enhance efficiency. By adjusting the distribution of argon flow, thermal load on the anode is reduced, which allows for significant energy savings. Experimental results show that a substantial portion of this saved energy is redirected to the engine’s power, thereby improving thrust efficiency. This radial segmentation method has led to efficiency values close to 60%, as energy that would have otherwise contributed to excess heating of the anode is instead utilized in propulsion. These findings suggest that, contrary to initial assumptions about argon’s lower efficiency, the strategic flow segmentation can make argon a viable and efficient choice for AF-MPDTs [16].
@@ 4 State of the Art: AF-MPDTs
Current AF-MPDT research and development is focused on the use of HTS coils to generate the magnetic field. Generally, previously built AF-MPDTs have had a magnetic field strength of up to 0.4 Teslas [40]. Creating a magnetic field with a magnetic field strength over 1 Tesla can theoretically maximize the performance of the thruster[20]. However, with conventional electromagnets, that strength of a magnetic field requires bulky designs with high power usage. The high current densities of superconductors enable the creation of lighter, more compact, electromagnets, allowing the thruster to avoid the heavy mass and power consumption tolls of the conventional electromagnet[41]. Multiple entities around the world are currently engaged with superconducting AF-MPDT research, including groups in Russia, Germany, China, and New Zealand[20]. The University of Stuttgart prototype currently is the only AF-MPDT design with a fully considered thermal management system; this remains a shortcoming of the Chinese, Russian, and New Zealand designs.
@@ 4.1 New Zealand - Robinson Research Institute (RRI)
4.2 Germany - University of Stuttgart Institute of Space Systems (IRS)
The Institute of Space Systems (IRS) is a research institute located at the University of Stuttgart in Germany. They have built a 100 kW AF-MPDT and worked with Neutron Star Systems, a German space startup, to build a 5 kW AF-MPDT, named SUPREME. Under the lead of the IRS, the SUPREME consortia is now developing both the 5 kW AF-MPDT and the nominal 100 kW AF-MPDT. The 5 kW AF-MPDT features a light and compact superconductor and cryocooler system. They are currently researching coil configurations, cryogenic technologies, and magnetic field shapes [20].
An important recent success comes from the Magnetohydrodynamic Enhanced Entry System for Space Transportation (MEESST) project, developed as part of a multi-institutional European constortium, including IRS, Neutron Star Systems, KU Leuven, and many other collaborators, which leverages HTS coils to generate strong magnetic fields that influence the plasma surrounding a spacecraft during atmospheric re-entry. This field pushes the ionized plasma layer away from the vehicle, significantly reducing the thermal load. By minimizing the need for traditional thermal protection systems, MEESST could lower spacecraft mass by up to 30%, a critical improvement for cost-effective space missions. Additionally, initial testing has shown MEESST’s HTS-equipped plasma probe achieving up to an 80% reduction in heat load, enabling a lighter and more compact approach to spacecraft shielding. By applying similar magnetic field configurations to AF-MPDTs, AF-MPDTs could achieve reduced thermal stresses on critical components, potentially increasing thrust efficiency and prolonging operational lifespan [43].
@@ 4.4 China - Chinese Academy of Sciences (CAS)
In 2022, the Chinese Academy of Sciences successfully designed and tested a 150 kW AF-MPDT with superconducting coils. They reported an efficiency of 76% and a specific impulse of over 5600 seconds. The thruster also demonstrated and confirmed the long-held theory that more powerful magnetic fields can reduce cathode erosion and offer a solution to AF-MPDT lifetime issues. However, it appears that this design did not incorporate a sufficient thermal management system. The lack of adequate cooling may have resulted in conditions where the cathode remained cooler than it would be in actual operating scenarios. In this cooler state, electron emission and plasma characteristics differ, resulting in data that likely underestimates the thermal stresses the system would encounter in actual operation. This discrepancy may lead to inaccurate assessments of the thruster’s performance and efficiency [40][20].
@@ Acknowledgements
The authors would like to thank CCIR for funding support and the opportunity to conduct research on this topic. HR is supported by the Schmidt Science Fellows. The authors also thank Georg Herdrich of IRS Stuttgart for his helpful comments and discussion.